OpenStudy (readergirl12):

Gene has a gasoline budget of $225 per month. He uses $5 of gasoline each day he drives. Which of the following equations represents how much money is left in his gasoline budget after x days of driving? A. y = 225x – 5 B. y = 225 – 5x C. y = 5x – 225 D. y = 225 + 5x

OpenStudy (texaschic101):

y = 225 - 5x

OpenStudy (gabebae):

@texaschic101 Explain to him/her

OpenStudy (texaschic101):

y = how much is left x = number of days 225 is what he started with....subtract 5 each day y = 225 - 5x so lets say he drove for 2 days....x = 2 y = 225 - 5(2) y = 225 - 10 y = 215 so in this case, he is left with $215 after driving 2 days

OpenStudy (readergirl12):

Thanks.Could you help me with some more? c:

OpenStudy (texaschic101):

sure

OpenStudy (readergirl12):

You mind if I tag you?

OpenStudy (texaschic101):

not at all...tag me if you need to